OpenStudy (drummmlers):

find the unknown value: 36^y = 6

rebeccaxhawaii (rebeccaxhawaii):

hint if 6^2=36 or 6*6=36 then what is 36^y=6?

OpenStudy (johnweldon1993):

As the lovely @rebeccaxhawaii has hinted at...you can rewrite this because you know \(\large 36 = 6^2\) So what you can write is \[\large (6^2)^y = 6^1\] *I just wrote the 1 exponent for later...so you can see* We know we can multiply these 2 exponents because \(\large (a^b)^c = a^{bc}\) so we have \(\large 6^{2y} = 6^1\) So now we merely equate these two...when does 6^(2y) = 6^(1) Well it happens when \(\large 2y = 1\) So what would \(\large y\) equal?

rebeccaxhawaii (rebeccaxhawaii):

smartcicles over here ^^